# Privacy Issues in Large Language Models: A Survey

### 23 Jan 2024 v2

@@ 2 General Terminology
Zhang et al. (2016)
Wikipedia
XML text dump of English Wikipedia
Mahoney (2009)
@@ 3 Memorization
If we define memorization broadly, all machine learning models will naturally exhibit memorization to some degree. Training a model necessitates observing training data and recalling that past information to make predictions. In the context of LLMs however, we want the models to generate content that semantically makes sense, while maintaining a degree of novelty. Initial studies on memorization have established that language models are capable of memorizing a significant fraction their training data, beginning with Carlini et al. (2019), who study LSTMs, and continuing with Carlini et al. (2020) who were able to extract $600$ memorized examples from GPT-2, and Carlini et al. (2023a) who study larger GPT-Neo models. Memorization is in itself, not necessarily an undesirable property of a machine learning model. In fact, recent work has show that a certain amount of memorization may be required for sufficiently accurate learning in certain natural classes of problems. Feldman (2020) first proposes a formal model to study this phenomenon, and shows that when the distribution of sub-populations in the training data is “long-tailed,” some amount of label memorization is required to achieve good generalization error. This has implications for next-token prediction on text data, which due to the variety and number of unique sentences may also be “long-tailed” in this sense. Brown et al. (2021) follow up on this work, and show that in similar natural settings the model may be forced to memorize training data that may even be extraneous to the prediction task, in order to generalize well. That being said, there are evident types of memorization that clearly pose a privacy risk given the presence of sensitive information in the training set, and serve no obvious purpose for generalization. Carlini et al. (2019) call this type of memorization “unintended memorization,” and show it occurs by inserting out of distribution samples called canaries during training, and demonstrating they can be extracted. Brown et al. (2021) formalizes this notion mathematically, by defining it as the mutual information between the generative model and the training data, conditional on the true distribution. For example, if a person named John Doe’s private information was present in the training data of a generative model, it could respond to the prompt “My name is John Doe and my phone number and address are…” with John Doe’s actual personal information Carlini et al. (2020). Or the model may output copyrighted data that is contained many times in the training set verbatim, for example the first act of Hamlet. More examples of these risks are discussed in Section 4 and Section 6 later on in this survey. On the other hand, some types of “memorization” are desired; for example when responding to the prompt “What is the capital city of the state of Rhode Island,” the desired output is of course “Providence.” Teasing apart these scenarios, and replacing intuition with mathematical definition is the content of the next subsection.
@@ 4 Privacy Attacks Against Language Models
In this section, we review recent works on privacy attacks against language models. This recent work builds upon a long line of research in the machine learning community that studies privacy attacks against ML models, typically in the classification setting. We refer the reader to Hu et al. (2022); Liu et al. (2021) for comprehensive surveys. Privacy attacks can largely be divided into two classes: membership inference attacks (MIAs), and training data extraction attacks (and attribute-inference attacks which we discuss in brief). MIAs are a less obviously severe privacy attack, where an adversary, given access to the model, is able to determine whether a given point was used to train that model with high accuracy. While this indicates that some information about the training point is “leaking” through the model, given that the adversary is still required to have access to a candidate point that may have been used to train the model in order to run the attack, it is more of a “smoke signal” than an outright privacy violation. Training data extraction on the other hand, is when given access to the model, an adversary can actually reconstruct parts of the underlying training data, which may be sensitive or protected by law.
@@ 4.2.1 Threshold Attacks
While the attack by Mireshghallah et al. (2022) was effective, it still relied on the assumption that the adversary has access to data that is very similar to the one used by the target. Mattern et al. (2023) devised a framework which does not rely on having access to any database. Instead, a separate model, called the proposal model, is trained to create “neighbors”, i.e. sequences that are very similar to the target sequence but with slight perturbations. These neighbors are fed into the target model and the difference between the target sequence and the average loss of the neighbor sequences is used at the threshold statistic (see Figure 9). This method was tested on a base version of GPT-2 Radford et al. (2019) trained on the AG News Corpus Gulli (2004), the Sentiment140 dataset Go (2009), and excerpts from the Wikitext-103 dataset Merity et al. (2016). Mattern et al. (2023) compared their method to two baselines, the loss only model by Yeom et al. (2018) and the likelihood ratio model by Mireshghallah et al. (2022) and found that it outperformed both. The TPR of Mattern et al. (2023) was marginally better or up to twice as high as the Mireshghallah et al. (2022) framework depending on the dataset. Both frameworks outperformed the baseline loss only model.
@@
Li et al. (2023) develop an MIA using a similar perturbation-based technique to Mattern et al. (2023), but rather than perturbing the sentence $x$, they perturb the model parameters themselves with Gaussian noise, and evaluate the resulting drop in the loss. This statistic corresponds to estimating the trace of the Hessian with respect to the model parameters around the $x$, via the Hutchinson Trace Estimator. They show that they are able to achieve improved attack accuracy relative to loss-based methods, including improved TPRs at low FPRs across the Pythia suite modelst trained on the Pile Biderman et al. (2023b). All methods, including loss-based ones and neighborhood comparison methods, fail to achieve non-trivial TPR at very low FPRs in this evaluation, even though MoPe makes the very strong assumption of white-box access to the model.
@@ 4.2.2 Shadow Model Attacks
Many ML models are built using a process known as transfer learning. In this, the model is pre-trained on a generic public dataset before being fine tuned on a case specific dataset. Abascal et al. (2023) implemented a shadow model MIA to infer membership in the pre-training dataset of a language model using only query access to the finetuned model. Since language models empirically forget over time Jagielski et al. (2023), inferring pre-training membership could be more challenging than inferring membership in the fine-tuning set. This attack was performed on the publicly available Transformer-XL foundation model Dai et al. (2019) which was pretrained on the WikiText-103 Merity et al. (2016) dataset and fine-tuned on the DBpedia dataset Zhang et al. (2016). The shadow models were built using these same public models and datasets and the attack model $f_{attack}$ (i.e. the model trained on the shadow model output) is a $k$-nearest neighbors model. As seen in Table 4, Abascal et al. (2023) was able to infer membership in pre-training datsets with TPRs as much as 10×10$\times$10 × higher than the baseline of randomly assigning membership. This attack shows the MIA risks are very much inherent for models which are publicly available, which removes the computational barrier to entry for a shadow model attack on an LLM. However, since pre-training datasets usually do not contain sensitive information, this is not as dangerous of a risk.
@@ 4.3 Training Data Extraction from Language Models
Carlini et al. (2020) improved upon this baseline in a few ways. On the candidate generation end, they tried two new methods. First, they used a decaying temperature setup (labeled Temperature in Table 5) where for early tokens, the confidence in top predictions was lowered to encourage diversity. As the generation continued, the “temperature” would decrease yielding more predictable generation. Second, they seeded the initial prefix using data from Common Crawl Crawl (2008) (Internet) to ensure there are no highly unusual prefixes. On the MIA approach end, instead of simply ranking by perplexity, Carlini et al. (2020) ranked by perplexity relative to a second, similar model or normalized by another value. Their comparison models include different sized GPT-2 models (Small and Medium) and a model trained on the lowercased version of the data (Lowercase). They also tried dividing the perplexity by the zlib entropy of the candidate (zlib) and also considering only the lowest perplexity from a sliding window of $50$ of the $256$ tokens. These methods ensure that common and uninteresting phrases are filtered out. As we see in Table 5, these improved methods enabled much more training data to be extracted. Of the $604$ total examples extracted, $46$ were named individuals not from a news source and $32$ were the contact information of individuals.
@@
This paper highlights that base language models of all sizes have a high propensity to regurgitate training data, and that alignment, at least the way state-of-the-art models are aligned, does not seem to solve this problem in the face of adversarial prompting strategies.
The rapid improvement in these attacks expose the significant risk of training data extraction in LLMs, with Patil et al. (2023) finding that there is no one universal solution to protecting against this risk. Similar work has also been done for image diffusion models Carlini et al. (2023c). While this is outside the scope of this survey, it highlights the prominence of this type of attack.
@@ 7.3 Unlearning in LLMs
Yao et al. (2023) similarly performed unlearning using gradient ascent, providing “negative” examples for the OPT and Llama models to unlearn. They showed that after several training iterations, these models were able to remove harmful content (such as recommended illegal activity), copyright content, and misinformation. Yao et al. (2023) compared this approach to reinforcement learning from human feedback (RLHF) and found gradient ascent unlearning achieved better goal alignment while using $2\%$ of the computational costs. This study showed that a gradient based approach to unlearning in LLMs can work in these practical cases.
@@
Concept Unlearning. While prior works focus on unlearning batches of examples, Eldan and Russinovich (2023) developed an approximate unlearning approach that can be used to unlearn an entire target data corpus, $D_{target}$. In this approach, a “reinforced” model is built by further fine tuning the model on $D_{target}$. The probability outputs of this model are compared to the baseline model since tokens common in $D_{target}$ will have higher probabilities than generic tokens. Generic counterpart labels are then generated and the baseline model is fine tuned using these alternate labels for several iterations until the target data is unlearned. Eldan and Russinovich (2023) tested this on Meta’s $7$-billion parameter Llama model and Microsoft’s $1.3$-billion parameter MSFT model and attempted to unlearn the Harry Potter books. After several fine-tuning iterations, the model lost its “memory” of Harry Potter. For example, the text completion prompt “Harry Potter’s two best friends are…” was completed as “a talking cat and a dragon. One day they decided to…” instead of “Ron Weasley and Hermione Granger. In this series…”. Model performance on other benchmarks decreased marginally but overall performance was largely unaffected.
@@
Zanella-Béguelin et al. (2020) found that model updates (such as unlearning) can possibly pose a privacy risk if an adversary has access to the versions of the model pre and post update. Through a differential analysis of the model output, a powerful membership inference attack can be performed on the model. For more details on this, see Section 4. Their attack highlights how machine unlearning is an orthogonal concept to differential privacy, and can in some cases increase risk from a data privacy perspective.

### 20 Feb 2024 v3

@@ 2 General Terminology
TLDR News
Daily tech newsletter collection
Belvèze (2022)
@@
Wikicorpus
Wikipedia in English, Spanish, and Catalan
Reese et al. (2010)
Wikipedia
XML text dump of English Wikipedia
Mahoney (2009)
@@
Zhang et al. (2016a)
XSum
Summaries of news articles
Narayan et al. (2018)
@@ 3 Memorization
If we define memorization broadly, all machine learning models will naturally exhibit memorization to some degree. Training a model necessitates observing training data and recalling that past information to make predictions. In the context of LLMs however, we want the models to generate content that semantically makes sense, while maintaining a degree of novelty. Initial studies on memorization have established that language models are capable of memorizing a significant fraction their training data, beginning with Carlini et al. (2019), who study LSTMs, and continuing with Carlini et al. (2020) who were able to extract $600$ memorized examples from GPT-2, and Carlini et al. (2023a) who study larger GPT-Neo models. Memorization is in itself, not necessarily an undesirable property of a machine learning model. In fact, recent work has show that a certain amount of memorization may be required for sufficiently accurate learning in certain natural classes of problems. Feldman (2020) first proposes a formal model to study this phenomenon, and shows that when the distribution of sub-populations in the training data is “long-tailed,” some amount of label memorization is required to achieve good generalization error. This has implications for next-token prediction on text data, which due to the variety and number of unique sentences may also be “long-tailed” in this sense. Brown et al. (2021) follow up on this work, and show that in similar natural settings the model may be forced to memorize training data that may even be extraneous to the prediction task, in order to generalize well. That being said, there are evident types of memorization that clearly pose a privacy risk given the presence of sensitive information in the training set, and serve no obvious purpose for generalization. Carlini et al. (2019) call this type of memorization “unintended memorization,” and show it occurs by inserting out of distribution samples called canaries during training, and demonstrating they can be extracted. Brown et al. (2021) formalizes this notion mathematically, by defining it as the mutual information between the generative model and the training data, conditional on the true distribution. For example, if a person named John Doe’s private information was present in the training data of a generative model, it could respond to the prompt “My name is John Doe and my phone number and address are…” with John Doe’s actual personal information Carlini et al. (2020). Or the model may output copyrighted data that is contained many times in the training set verbatim, for example the first act of Hamlet. More examples of these risks are discussed in Section 4 and Section 6 later on in this survey. On the other hand, some types of “memorization” are desired; for example when responding to the prompt “What is the capital city of the state of Rhode Island,” the desired output is, of course, “Providence.” Teasing apart these scenarios, and replacing intuition with mathematical definition is the content of the next subsection.
@@ 3.3 Mitigation Techniques
LLM Editing: Another approach to reducing memorization is to directly edit the neurons of the LLM that correspond to the memorization. Chang et al. (2023) tested several localization methods that identify the set of LLM parameters responsible for storing certain knowledge. They found that their method known as Hard Concrete was able to identify sets of neurons consisting of approximately .$5\%.$ of the Pythia $6.9$B model where removing the neurons reduced the memorization of that example by over $50\%$. Other localization methods were able to achieve varying degrees of success but across all of them, targeted memorization reduction also resulted in a drop in performance. Despite this, the approach of editing LLMs to protect sensitive information is a promising direction for memorization mitigation.
Mitigation For Translation Models: Neural machine translation (NMT) models are another application of LLMs for the purposes of language translation. Within this space, Raunak and Menezes (2022) adapted the extractability notion (Definition 1) of memorization to NMT models and showed that in addition to memorization occurring in NMT models, many neighbors of the inputs (i.e. inputs that differ in one word) output the same memorized results. Their proposed algorithm appends the input with a special character (e.g. "!"), then removes the character after translation. This approach results in $65\%$ of memorized translations returning a new, but still semantically correct, output.
@@ 4 Privacy Attacks Against Language Models
In this section, we review recent works on privacy attacks against language models. This recent work builds upon a long line of research in the machine learning community that studies privacy attacks against ML models, typically in the classification setting. We refer the reader to Hu et al. (2022) and Liu et al. (2021) for comprehensive surveys. Privacy attacks can largely be divided into two classes: membership inference attacks (MIAs), and training data extraction attacks (and attribute-inference attacks which we discuss in brief). MIAs are a less obviously severe privacy attack, where an adversary, given access to the model, is able to determine whether a given point was used to train that model with high accuracy. While this indicates that some information about the training point is “leaking” through the model, given that the adversary is still required to have access to a candidate point that may have been used to train the model in order to run the attack, it is more of a “smoke signal” than an outright privacy violation. Training data extraction on the other hand, is when given access to the model, an adversary can actually reconstruct parts of the underlying training data, which may be sensitive or protected by law.
@@ 4.2.1 Threshold Attacks
While the attack by Mireshghallah et al. (2022) was effective, it still relied on the assumption that the adversary has access to data that is very similar to the one used by the target. Mattern et al. (2023) devised a framework which does not rely on having access to any database. Instead, a separate model, called the proposal model, is trained to create “neighbors”, i.e. sequences that are very similar to the target sequence but with slight perturbations. These neighbors are fed into the target model and the difference between the target sequence and the average loss of the neighbor sequences is used at the threshold statistic (see Figure 9). This method was tested on a base version of GPT-2 Radford et al. (2019) trained on the AG News Corpus Gulli (2004), the Sentiment140 dataset Go (2009), and excerpts from the Wikitext-103 dataset Merity et al. (2016). Mattern et al. (2023) compared their method to two baselines, the loss only model by Yeom et al. (2018) and the likelihood ratio model by Mireshghallah et al. (2022) and found that it outperformed both. The TPR of Mattern et al. (2023) was at least marginally better and up to twice as high as the Mireshghallah et al. (2022) framework depending on the dataset. Both frameworks outperformed the baseline loss only model.
@@
Fu et al. (2023) developed a MIA, known as Self-calibrated Probabilistic Variation (SPV-MIA) that shares the similar perturbation approach but with a few notable differences. As seen in Figure 10, the variation in performance of the target model on the perturbed neighbors is compared to a reference model rather than a strict threshold. This was designed so that the success of the attack is not dependent on model overfitting, which Fu et al. (2023) argues will soon be mitigated by regularization methods and improved generalization of LLMs. Reference model-based attacks rely on the ability for the reference dataset to closely resemble the target dataset $D_{train}$. Fu et al. (2023) resolves this by fine-tuning the reference model using a dataset generated by prompting the target model. Due to the self-prompting nature of this fine-tuning, this attack can be performed without prior knowledge of $D_{train}$. They tested this attack on GPT-2, GPT-J, Falcon-7B, and LLaMA-7B models using the Wikitext-$103$, AG News, and XSum datasets for training with Wikicorpus, TLDR News, and CNNDM datasets used as the initial datasets for the respective reference models. They compared this attack against other MIAs such as the threshold attack by Yeom et al. (2018), LiRA attacks, and Mattern et al. (2023) and saw a marked improvement in AUC of up to $30\%$ with AUC on all datasets and models hovering at or above $90\%$.
Figure 10: Framework of SPV-MIA Fu et al. (2023). A paraphrasing model generates neighbors of the target. These examples are fed to both the target model and a reference model (fine-tuned using output from the target LLM) and the difference in the probabilistic variations determines the membership threshold.
Li et al. (2023) also developed an MIA using a perturbation-based technique, but rather than perturbing the sentence $x$, they perturb the model parameters themselves with Gaussian noise, and evaluate the resulting drop in the loss. This statistic corresponds to estimating the trace of the Hessian with respect to the model parameters around the $x$, via the Hutchinson Trace Estimator. They show that they are able to achieve improved attack accuracy relative to loss-based methods, including improved TPRs at low FPRs across the Pythia suite modelst trained on the Pile Biderman et al. (2023b). All methods, including loss-based ones and neighborhood comparison methods, fail to achieve non-trivial TPR at very low FPRs in this evaluation, even though MoPe makes the very strong assumption of white-box access to the model.
@@ 4.2.2 Shadow Model Attacks
Many ML models are built using a process known as transfer learning. In this, the model is pre-trained on a generic public dataset before being fine tuned on a case specific dataset. Abascal et al. (2023) implemented a shadow model MIA to infer membership in the pre-training dataset of a language model using only query access to the finetuned model. Since language models empirically forget over time Jagielski et al. (2023), inferring pre-training membership could be more challenging than inferring membership in the fine-tuning set. This attack was performed on the publicly available Transformer-XL foundation model Dai et al. (2019) which was pretrained on the WikiText-103 Merity et al. (2016) dataset and fine-tuned on the DBpedia dataset Zhang et al. (2016a). The shadow models were built using these same public models and datasets and the attack model $f_{attack}$ (i.e. the model trained on the shadow model output) is a $k$-nearest neighbors model. As seen in Table 4, Abascal et al. (2023) was able to infer membership in pre-training datsets with TPRs as much as 10×10$\times$10 × higher than the baseline of randomly assigning membership. This attack shows the MIA risks are very much inherent for models which are publicly available, which removes the computational barrier to entry for a shadow model attack on an LLM. However, since pre-training datasets usually do not contain sensitive information, this is not as dangerous of a risk.
@@ 4.3 Training Data Extraction from Language Models
Carlini et al. (2020) improved upon their baseline in a few ways. On the candidate generation end, they tried two new methods. First, they used a decaying temperature setup (labeled Temperature in Table 5) where for early tokens, the confidence in top predictions was lowered to encourage diversity. As the generation continued, the “temperature” would decrease yielding more predictable generation. Second, they seeded the initial prefix using data from Common Crawl Crawl (2008) (Internet) to ensure there are no highly unusual prefixes. On the MIA approach end, instead of simply ranking by perplexity, Carlini et al. (2020) ranked by perplexity relative to a second, similar model or normalized by another value. Their comparison models include different sized GPT-2 models (Small and Medium) and a model trained on the lowercased version of the data (Lowercase). They also tried dividing the perplexity by the zlib entropy of the candidate (zlib) and also considering only the lowest perplexity from a sliding window of $50$ of the $256$ tokens. These methods ensure that common and uninteresting phrases are filtered out. As we see in Table 5, these improved methods enabled much more training data to be extracted. Of the $604$ total examples extracted, $46$ were named individuals not from a news source and $32$ were the contact information of individuals.
@@
This paper highlights that base language models of all sizes have a high propensity to regurgitate training data, and that alignment, at least the way state-of-the-art models are aligned, does not seem to solve this problem in the face of adversarial prompting strategies. The widespread use of these LLMs such as ChatGPT or OpenAI’s fine-tuning API lead to expanded risks of sensitive data leaks Sun et al. (2023).
The rapid improvement in these attacks expose the significant risk of training data extraction in LLMs, with Patil et al. (2023) finding that there is no one universal solution to protecting against this risk. This risk also extends beyond natural language use cases alone. Al-Kaswan et al. (2023) found that code generation models such as CodeGen, CodeParrot, and PyCodeGPT are susceptible to training data extraction and were able to extract up to $47\%$ of the training code. Further extraction attacks have been done for image diffusion models Carlini et al. (2023c). While this is outside the scope of this survey, it highlights the prominence of this type of attack.
@@ 5.2.2 DP Fine-Tuning of LLMs
Similar to Li et al. (2022), Yu et al. (2022) looked to use DP fine-tuning on much large LLMs but improved computational constraints by limiting the parameters that are updated in fine-tuning. As shown in Figure 18, Yu et al. (2022) starts with a model that was pre-trained on a public dataset. They then add additional parameters to the model (where the number of new parameters is much smaller than in the original model) and fine-tune using DP-SGD with the private dataset. Importantly, Yu et al. (2022) freezes the parameters from the pre-trained model. By doing this, they reduce computational costs of fine-tuning while also reducing the noise added to the model. This approach was tested starting with RoBERTa models Liu et al. (2019), which are models pre-trained on public internet data. They were fine-tuned to attempt the MNLI, QQP, SST-2, and QNLI tasks from the GLUE benchmark Wang et al. (2018). Using a privacy budget of $\epsilon=6.7$, they were able to achieve an average accuracy of $90.3\%$ across these four tasks (non-private fine-tuning achieved an average accuracy of $93.4\%$), showing that this approach is effective at achieving privacy with only modest drop in performance. An added benefit to this framework is its flexibility in multiple use cases. For each new task, a new set of fine-tuned parameters can be trained and then added on to the base model, reducing the need to have multiple copies of the large base model.
Figure 18: Illustration of the private fine-tuning framework from Yu et al. (2022). In this approach, new parameters are added to a pre-trained model and are fine-tuned on the private dataset. The original parameters are frozen during this process. The fine-tuned parameters can then be released publicly and plugged into the model for future use cases, while maintaining the privacy of the private dataset.
@@ 7.3 Unlearning in LLMs
Yao et al. (2023) similarly performed unlearning using gradient ascent, providing “negative” examples for OPT and Llama models to unlearn. They showed that after several training iterations, these models were able to remove harmful content (such as recommended illegal activity), copyright content, and misinformation. Yao et al. (2023) compared this approach to reinforcement learning from human feedback (RLHF) and found gradient ascent unlearning achieved better goal alignment while using $2\%$ of the computational costs. This study showed that a gradient based approach to unlearning in LLMs can work in these practical cases.
Kassem et al. (2023) developed a method for unlearning called DeMem (standing for DeMemorization) which takes a gradient based solution using a reinforcement learning feedback loop. Given a dataset that is split into prefixes and suffixes (e.g. "Alice lives at…" and "…123 main street"), DeMem uses an independent, pre-trained LM to produce a new suffix. Then a similarity score is computed between this suffix and the original one. This score is used as the reward signal in the feedback loop, encouraging the target model to develop a paraphrasing policy. Kassem et al. (2023) tested this method on GPT-Neo and OPT models that were trained on the Pile. BERTScore was used as the similarity score Zhang et al. (2020), memorization was measured by edit distance using SacreBLEU Post (2018), and model performance was measured across $9$ different task benchmarks and $2$ perplexity benchmarks. Combining DeMem with deduplication Kandpal et al. (2022) saw a marked decrease in memorized examples, between half and a third as many as the baseline. Meanwhile, performance only marginally decreased by about .$5\%.$ across the benchmarks. The robustness of the performance comes from the use of the pre-trained LM which ensures that the updated model remains coherent.
In-Context Unlearning: While the above methodologies are proven unlearning approaches, they all require access to the model parameters in some capacity in order to perform the unlearning updates. Pawelczyk et al. (2023) provides a new framework for unlearning they call In-Context Unlearning (ICUL). This takes inspiration from the concept of In-Context Learning (ICL) where examples are provided to the LLM as part of the prompt in order to improve model performance for a specific task without having to update model parameters. Given a deletion request for a certain training example, ICUL provides that instance with a flipped label along with several additional, correctly labeled instances which are prepended to the input to the LLM. This approach was tested on the $560$M and $1.1$B parameter Bloom language models Workshop et al. (2023) with the SST-2 Socher et al. (2013), Amazon polarity and Yelp polarity Zhang et al. (2016b) datasets. These datasets are text classification tasks with the goal of predicting positive or negative reviews on Rotten Tomatoes, Amazon, and Yelp respectively. The performance of ICUL was compared to two baselines, one being ICL which does not provide the flipped examples and the other being the unlearning approach by Jang et al. (2023). These approaches were all measured by their performance against a LiRA attack using $10$ shadow models Carlini et al. (2022).
ICUL surpassed the baseline with lower AUC and TPR (meaning the LiRA attack was less successful) across both model sizes and all datasets. Compared to Jang et al. (2023), ICUL achieved lower AUC scores in all instances and lower TPR in the majority of cases. These results show that ICUL greatly reduced the ability to identify the forgotten point as a member of the training set with the added advantage of only having black-box access to the model. ICUL achieved AUC scores that were near .5.5.5.5 meaning that these results are non-trivial in a practical setting. While task performance suffers slightly with this approach, this difference shrinks on the larger Bloom model with test accuracy being within $1\%$ of the baseline model on the polarity datasets.
Two additional studies were run on this approach to determine its effectiveness. First, Pawelczyk et al. (2023) varied the number of correct samples provided after the flipped example to determine the dependency on these. Pawelczyk et al. (2023) found that short context lengths (i.e. $2$ additional examples) yielded poor model confidence scores. Longer context lengths of $4$ or $6$ yielded good performance. Second, they tested the dependency on the forget instance by keeping the flipped label but replacing the training input with a randomly sampled input. ICUL significantly outperforms this randomized approach showing that the flipped label step is essential in this unlearning process.
Figure 26: An example of the In-Context Unlearning approach compared to in-context learning from Pawelczyk et al. (2023). For the given sentiment analysis task, in-context learning provides labeled examples from the training set to help the LLM answer the prompt. In-context unlearning flips the label of the forget point (“Over and over again.”) before providing correctly labeled examples. This process reduces the ability of a LiRA attack to identify members of the training set.
@@
Concept Unlearning: While prior works focus on unlearning batches of examples, Eldan and Russinovich (2023) developed an approximate unlearning approach that can be used to unlearn an entire target data corpus, $D_{target}$. In this approach, a “reinforced” model is built by further fine tuning the model on $D_{target}$. The probability outputs of this model are compared to the baseline model since tokens common in $D_{target}$ will have higher probabilities than generic tokens. Generic counterpart labels are then generated and the baseline model is fine tuned using these alternate labels for several iterations until the target data is unlearned. Eldan and Russinovich (2023) tested this on Meta’s $7$-billion parameter Llama model and Microsoft’s $1.3$-billion parameter MSFT model and attempted to unlearn the Harry Potter books. After several fine-tuning iterations, the model lost its “memory” of Harry Potter. For example, the text completion prompt “Harry Potter’s two best friends are…” was completed as “a talking cat and a dragon. One day they decided to…” instead of “Ron Weasley and Hermione Granger. In this series…”. Model performance on other benchmarks decreased marginally but overall performance was largely unaffected.
Figure 27: Shi et al. (2023), Auditing machine unlearning with Min-KProb. Machine unlearning methods are designed to remove copyrighted and personal data from large language models. Min-Kis used to audit an unlearned LLM that has been trained to forget copyrighted books. However, such a model can still output related copyrighted content
@@
Zanella-Béguelin et al. (2020) found that model updates (such as unlearning) can possibly pose a privacy risk if an adversary has access to the versions of the model pre and post update. Through a differential analysis of the model output, a powerful membership inference attack can be performed on the model. For more details on this, see Section 4. Their attack highlights how machine unlearning is not an orthogonal concept to differential privacy, and can in some cases increase risk from a data privacy perspective.